A ball is thrown at an initial height of 7.5 feet with an initial velocity of 130 feet per second. The following equation gives the height h(t) of the ball in feet after t seconds.
h(t) = -16t^2 + 130t + 7.5
How long will it take the ball to hit a window at a height of 30 feet assuming it strikes on the downward path?

Answer by TimothyLamb(4379)

---
h(t) = -16t^2 + 130t + 7.5 = 30
-16t^2 + 130t - 22.5 = 0
---
the above quadratic equation is in standard form, with a=-16, b=130, and c=-22.5
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
-16 130 -22.5
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the two real roots (i.e. the two x-intercepts), of the quadratic are:
t = 0.176929739
t = 7.94807026
---
answer:
the ball is at 30 ft and falling at time t=7.948 seconds.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php